# A joint diffusion approach to multi-modal inference in inertial confinement fusion

### v1

@@ Introduction
Inertial confinement fusion (ICF) has emerged as a leading approach for controlled nuclear fusion (?, ?), providing fundamental plasma physics insights and a potential pathway toward alternative energy. Scientific progress in ICF relies on a close interplay between simulation and experiment, in which each is iteratively refined based on data from the other. Machine learning models can enhance this interplay by acting as fast surrogates for simulations and by identifying input conditions that reproduce observed experimental outputs (?, ?, ?, ?). In practice, however, each implosion experiment yields a distinct and limited set of diagnostics, whereas simulations produce a richer and more uniform set of outputs. Moreover, these outputs are multi-modal – a combination of scalars and images – making it challenging to learn flexible conditional distributions based only on the available subset of diagnostics (?).
@@
Multi-modal generative models have become powerful tools across scientific domains, enabling advances in biomolecular structure prediction, medical diagnostics, and materials design (?, ?, ?). These models capture complex relationships across heterogeneous data types and often employ generative decoders such as denoising diffusion probabilistic models (DDPMs) (?, ?) to generate high-quality samples. Multi-modal prediction is typically achieved either by combining data into a shared latent space prior to diffusion (?, ?) or by aligning pre-existing latent representations using contrastive learning techniques (?). While contrastive approaches allow the reuse of pre-trained models without exhaustive retraining, they may fail to capture subtle inter-modal dependencies by not explicitly modeling the full joint distribution (?).
@@
An alternative strategy is to train a generative model directly on the joint distribution of multi-modal data. This approach has gained traction in domains such as generative biology, where diffusion models jointly update continuous atom positions and discrete atom types (?, ?, ?). It has further been shown that joint diffusion is feasible and theoretically well-founded in arbitrary state spaces (?, ?), however the application of joint diffusion models as physics-based surrogates have been under-explored.
@@
We propose JointDiff, a model architecture and training scheme that leverages joint diffusion as a generative surrogate model for ICF. We train our model on a large ensemble of 3D Multi-Rocket Piston (RP) simulations (?) and demonstrate accuracy in forward and inverse modeling tasks as well as imputation of missing outputs. We show that JointDiff produces meaningful conditional distributions from partial multi-modal observations, enabling tunable uncertainty quantification and investigation of input sensitivity to missing diagnostics, with demonstrated transferability to experiments at the National Ignition Facility (NIF).
@@
Deep learning models have previously been applied as multi-modal ICF surrogates (?, ?) and techniques such as Markov Chain Monte Marlo (MCMC) have been used to reconstruct simulations inputs from experimental outputs (?, ?, ?), however implementing MCMC with multi-modal data is sensitive to choice of priors, computationally intensive, and may fail to converge. DDPMs have been applied as surrogates for Particle-in-cell (PIC) simulations (?), which are crucial to understand fundamental plasma interactions during ICF implosions, and preliminary work showed that DDPMs can model 2D radiation hydrodynamics simulations outputs with single-channel images (?). In this work, we generalize this framework to 3D simulation data using an multi-modal U-net based architecture amenable to arbitrary image views, images channels, and scalars inputs and outputs (Figure 2).
@@ Results
To predict conditional distributions given multi-modal (scalar and image) data, we use a joint diffusion objective (?, ?). During training, noise is gradually added to both scalars and images and a single architecture learns to “denoise” both modalities at the same time. We apply random masking over inputs and outputs such that the models learns the distribution of outputs given inputs (forward model) as well of inputs given outputs (inverse model). Output masks are varied individually over scalars and images, enabling prediction of inputs and missing outputs conditioned on partial observations. When making predictions, the model is guided by whatever data is available, and the prediction targets are masked. The model architecture is shown in Figure 2 and additional details are provided in Methods.
We train our model on over 443 thousand simulations generated by a RP physics model. The RP model uses a discretized shell to describe how fuel in a target capsule is compressed to form a hotspot; more detail can be found in Methods and in (?, ?). Each simulation training example contains a set of 28 scalar inputs, 12 output scalars, and 3 image-line-of-sight each with 2 energy channels. The input scalars represent parameters such as initial pressure, adiabat, and drive symmetry modes that define each simulation. Scalar outputs include quantities that can be directly compared to experiment, such as the total neutrons generated by the implosion (Yield) and the time of peak neutron production (Bang-time), as well as inferred quantities that cannot be measured experimentally such as the areal density (R) and residual kinetic energy (RKE), which are measures of hot spot confinement and energy not coupled to the hot spot, respectively. The images correspond to three different views of the implosion, with each image containing two channels: one for higher-energy (Primary) neutrons and one for lower-energy (Down-scattered) neutrons. A complete list of inputs and outputs can be found in the Supplementary Text.
@@
Next we test the inverse model’s ability to adapt and predict inputs when a larger fraction of output data is unavailable. In this test we specifically remove outputs that are regularly unavailable in NIF shots. For example, in NIF experiments View 3 is often blocked by other diagnostics and the View 2 (polar) detector cannot capture down-scattered neutrons as the camera’s line of sight is not long enough to separate neutron energies in time (?). Accordingly, we mask both channels of View 3 as well as the down-scatter channel of View 2. Furthermore, 4/12 scalars in our simulation dataset, measuring R and RKE quantities, are inferred outputs that have no direct experimental equivalent. Burn-width measurements can also be challenging at high yield where burn duration is comparable to detector resolution, and the RP Bang-time tends to be systematically lower than experiments (?). Therefore we remove half (6/12) of scalar outputs and half (3/6) of image channels, and we test the model’s ability to predict meaningful capsule inputs distribution. We emphasize that this model was not re-trained or fine-tuned for this specific mask configuration and it therefore retains the forward and inverse modeling capabilities showcased above.
@@
We perform the same round-trip analysis on images. Images channels provided to both models are shown in the green box in Figure 5C, and masked images are shown in the red box. Again we observe strong round-trip reconstructions when comparing the mean round-trip prediction to the original images, although there are some features slightly blurred in the partial round-trips. Taken together, these results verify that the predicted inputs distributions correctly correspond to their conditioning values and that the forward and inverse modes are self-consistent, even when given partial information and without strict self-consistency enforced as in (?).
@@
In Figure 6B we examine the round-trip distribution of scalars and images predicted for the recent N230729 and N240907 shots. We observe that predicted outputs distributions are wider than simulations, indicating higher uncertainty, but that ground-truth outputs are captured within the distributions. Across all shots, DSR is the hardest scalar to match, potentially indicating inconsistency in the underlying RP model. In Figure 6C we show round-trip predictions of View 1 primary neutron images for all experiments. Overall image shape is well captured in reconstructions, however more subtle features are blurred or lost, especially in N221204 and N240210. These features may not be present in the RP training data, and fine-tuning JointDiff on experiments or higher fidelity simulations (?, ?) would likely improve round-trip self-consistency.
@@ Discussion
Looking ahead, several directions may further improve model performance and generalizability. For ICF applications, we plan to fine-tune JointDiff using higher-fidelity simulation data from codes such as HYDRA (?), as well as experimental data. For more complex datasets, future work could incorporate the joint diffusion objective into advanced architectures like the diffusion transformer (?, ?), apply curriculum learning strategies (?) to incrementally introduce challenging mask configurations, or project images into pre-trained latent spaces before diffusion (?). Finally, integrating experimental uncertainties during inference may help constrain predicted distributions and better align model outputs with results from MCMC approaches.
@@ Acknowledgments
Supplementary Materials for
A joint diffusion approach to multi-modal inference in inertial confinement fusion
Michael Jones∗, Justin Kunimune, Daniel Casey, Bogdan Kustowski, Eugene Kur, Kelli Humbird ∗Corresponding author. Email: jones313@llnl.gov
@@
Simulated implosions are modeled using a radiation drive derived from the Callahan hohlraum model (?) and solve the rocket equations for 200 coupled rocket-pistons that discretize the capsule shell in 3D. After the initial acceleration phase, the pistons are linked through hotspot pressure using power-balance equations, following the framework of Springer et al. (?) but with an explicit hohlraum model applied from the onset of implosion. Synthetic diagnostics at stagnation and peak burn are generated through post-processing. Full methodological details are provided in (?).
@@
An ensemble of 443,610 RP simulations was used with inputs sampled near the predicted experimental ignition cliff (?). A random split of 986 samples (0.2%) was held out for testing, and 100 additional samples were used for validation during training. Images were normalized to by dividing by max pixel intensity for each view and channel. Scalar inputs and outputs were normalized to given mean and then passed through an inverse sigmoid to more closely approximate zero-centered Gaussians. All evaluation metrics reported in Tables 1 and 2 and Figure 3 are performed after removing normalization.
@@
Deuterium-Tritium ICF experiments were performed at the National Ignition Facility (NIF) between June 2021 and September 2024. Additional information on specific shots can be found in (?, ?, ?, ?, ?, ?).
@@
DDPMs generate high resolution data distributions by first noising training data to gradually approach a Gaussian distribution, then iteratively denoising samples drawn from a Gaussian prior (?, ?). During training, samples can be drawn at arbitrary diffusion times according to , where , and corresponding to the noise schedule. We use a linear noise schedule that decreases from 0.9999 to 0.98 over 1000 steps.
@@
A multi-modal neural network architecture is used to encode and predict noise in both scalar and image pixel space. We adapt a U-net architecture (?) developed by (?) to ingest and predict multi-modal data. Each image view is embedded by independent convolutions networks containing 48 channels and 3 up/down sampling layers. Both input scalars, output scalars, and mask tokens are embedded by fully connected neural networks containing one 64-dimensional hidden layers. Diffusion time is embedded by a sin/cos positional encoding. Scalar, mask, and time embeddings are concatenated and added to U-net embedding at each up and down sampling step. Image embedding for each view receive the same time, scalar, and mask conditioning but are independent from each other until the bottleneck layer (lowest dimension of the U-net). At the bottleneck layer outputs from each image view are pooled along the channel dimension, flattened, and concatenated along with scalar and mask embedding. Output and input scalar decoders take the flattened bottleneck layer as input and predict noise for each scalar. Image decoders pool information from the bottleneck layer and upsample three times to recover the original pixel and channel dimensions.
Models are constructed in PyTorch (?), and trained using the Adam optimizer (?). We use a batch size of 256 and learning rate of 0.0003. An exponential moving average with constant 0.995 is applied to weights. Additional architecture details and hyper-parameters are included in Tables S1 and S2).
@@
Binary masks inform the model if a conditioning variable (scalar or image) is provided for conditioning, or if it should be predicted during the denoising (inference) process. Certain masks are linked to others if those quantities can never be known without the other. Specifically, all inputs masks are linked since they must all be specified to perform a simulation and cannot be partially observed in experiment. The four inferred outputs scalars (R Weighted Harmonic Mean, R Mean, RKE at Bang-time, and RKE at Minimum Volume) are linked for the same reason (?). The three velocity scalars are linked because they are x,y,z components of the same vector. All other output scalars have individual masks along with each image view and channel. During training, all independent masks are sampled according to a Bernoulli distribution with constant 0.5. Ablations were performed with varied constant values and with curriculum learning strategies to gradually lower the constant as function of training time, but a fixed constant of 0.5 provided optimal performance across forward, inverse, and imputation tasks.

### v2

@@ Introduction
Inertial confinement fusion (ICF) has emerged as a leading approach for controlled nuclear fusion 41, 3, providing fundamental plasma physics insights and a potential pathway toward alternative energy. Scientific progress in ICF relies on a close interplay between simulation and experiment, in which each is iteratively refined based on data from the other. Machine learning models can enhance this interplay by acting as fast surrogates for simulations and by identifying input conditions that reproduce observed experimental outputs 4, 25, 14, 37. In practice, however, each implosion experiment yields a distinct and limited set of diagnostics, whereas simulations produce a richer and more uniform set of outputs. Moreover, these outputs are multi-modal – a combination of scalars and images – making it challenging to learn flexible conditional distributions based only on the available subset of diagnostics 24.
@@
Multi-modal generative models have become powerful tools across scientific domains, enabling advances in biomolecular structure prediction, medical diagnostics, and materials design 1, 40, 39. These models capture complex relationships across heterogeneous data types and often employ generative decoders such as denoising diffusion probabilistic models (DDPMs) 15, 36 to generate high-quality samples. Multi-modal prediction is typically achieved either by combining data into a shared latent space prior to diffusion 5, 26 or by aligning pre-existing latent representations using contrastive learning techniques 32. While contrastive approaches allow the reuse of pre-trained models without exhaustive retraining, they may fail to capture subtle inter-modal dependencies by not explicitly modeling the full joint distribution 10.
@@
An alternative strategy is to train a generative model directly on the joint distribution of multi-modal data. This approach has gained traction in domains such as generative biology, where diffusion models jointly update continuous atom positions and discrete atom types 11, 31, 8. It has further been shown that joint diffusion is feasible and theoretically well-founded in arbitrary state spaces 16, 33, however the application of joint diffusion models as physics-based surrogates have been under-explored.
@@
We propose JointDiff, a model architecture and training scheme that leverages joint diffusion as a generative surrogate model for ICF. We train our model on a large ensemble of 3D Multi-Rocket Piston (RP) simulations 9 and demonstrate accuracy in forward and inverse modeling tasks as well as imputation of missing outputs. We show that JointDiff produces meaningful conditional distributions from partial multi-modal observations, enabling tunable uncertainty quantification and investigation of input sensitivity to missing diagnostics, with demonstrated transferability to experiments at the National Ignition Facility (NIF).
@@
Deep learning models have previously been applied as multi-modal ICF surrogates 4, 25 and techniques such as Markov Chain Monte Marlo (MCMC) have been used to reconstruct simulations inputs from experimental outputs 14, 24, 37, however implementing MCMC with multi-modal data is sensitive to choice of priors, computationally intensive, and may fail to converge. DDPMs have been applied as surrogates for Particle-in-cell (PIC) simulations 27, which are crucial to understand fundamental plasma interactions during ICF implosions, and preliminary work showed that DDPMs can model 2D radiation hydrodynamics simulations outputs with single-channel images 21. In this work, we generalize this framework to 3D simulation data using an multi-modal U-net based architecture amenable to arbitrary image views, images channels, and scalars inputs and outputs (Figure 2).
@@ Results
To predict conditional distributions given multi-modal (scalar and image) data, we use a joint diffusion objective 16, 32. During training, noise is gradually added to both scalars and images and a single architecture learns to “denoise” both modalities at the same time. We apply random masking over inputs and outputs such that the models learns the distribution of outputs given inputs (forward model) as well of inputs given outputs (inverse model). Output masks are varied individually over scalars and images, enabling prediction of inputs and missing outputs conditioned on partial observations. When making predictions, the model is guided by whatever data is available, and the prediction targets are masked. The model architecture is shown in Figure 2 and additional details are provided in Methods.
We train our model on over 443 thousand simulations generated by a RP physics model. The RP model uses a discretized shell to describe how fuel in a target capsule is compressed to form a hotspot; more detail can be found in Methods and in 9, 24. Each simulation training example contains a set of 28 scalar inputs, 12 output scalars, and 3 image-line-of-sight each with 2 energy channels. The input scalars represent parameters such as initial pressure, adiabat, and drive symmetry modes that define each simulation. Scalar outputs include quantities that can be directly compared to experiment, such as the total neutrons generated by the implosion (Yield) and the time of peak neutron production (Bang-time), as well as inferred quantities that cannot be measured experimentally such as the areal density (R) and residual kinetic energy (RKE), which are measures of hot spot confinement and energy not coupled to the hot spot, respectively. The images correspond to three different views of the implosion, with each image containing two channels: one for higher-energy (Primary) neutrons and one for lower-energy (Down-scattered) neutrons. A complete list of inputs and outputs can be found in the Supplementary Text.
@@
Next we test the inverse model’s ability to adapt and predict inputs when a larger fraction of output data is unavailable. In this test we specifically remove outputs that are regularly unavailable in NIF shots. For example, in NIF experiments View 3 is often blocked by other diagnostics and the View 2 (polar) detector cannot capture down-scattered neutrons as the camera’s line of sight is not long enough to separate neutron energies in time 13. Accordingly, we mask both channels of View 3 as well as the down-scatter channel of View 2. Furthermore, 4/12 scalars in our simulation dataset, measuring R and RKE quantities, are inferred outputs that have no direct experimental equivalent. Burn-width measurements can also be challenging at high yield where burn duration is comparable to detector resolution, and the RP Bang-time tends to be systematically lower than experiments 9. Therefore we remove half (6/12) of scalar outputs and half (3/6) of image channels, and we test the model’s ability to predict meaningful capsule inputs distribution. We emphasize that this model was not re-trained or fine-tuned for this specific mask configuration and it therefore retains the forward and inverse modeling capabilities showcased above.
@@
We perform the same round-trip analysis on images. Images channels provided to both models are shown in the green box in Figure 5C, and masked images are shown in the red box. Again we observe strong round-trip reconstructions when comparing the mean round-trip prediction to the original images, although there are some features slightly blurred in the partial round-trips. Taken together, these results verify that the predicted inputs distributions correctly correspond to their conditioning values and that the forward and inverse modes are self-consistent, even when given partial information and without strict self-consistency enforced as in 4.
@@
In Figure 6B we examine the round-trip distribution of scalars and images predicted for the recent N230729 and N240907 shots. We observe that predicted outputs distributions are wider than simulations, indicating higher uncertainty, but that ground-truth outputs are captured within the distributions. Across all shots, DSR is the hardest scalar to match, potentially indicating inconsistency in the underlying RP model. In Figure 6C we show round-trip predictions of View 1 primary neutron images for all experiments. Overall image shape is well captured in reconstructions, however more subtle features are blurred or lost, especially in N221204 and N240210. These features may not be present in the RP training data, and fine-tuning JointDiff on experiments or higher fidelity simulations 17, 25 would likely improve round-trip self-consistency.
@@ Discussion
Looking ahead, several directions may further improve model performance and generalizability. For ICF applications, we plan to fine-tune JointDiff using higher-fidelity simulation data from codes such as HYDRA 28, as well as experimental data. For more complex datasets, future work could incorporate the joint diffusion objective into advanced architectures like the diffusion transformer 30, 33, apply curriculum learning strategies 6 to incrementally introduce challenging mask configurations, or project images into pre-trained latent spaces before diffusion 34. Finally, integrating experimental uncertainties during inference may help constrain predicted distributions and better align model outputs with results from MCMC approaches.
@@ Acknowledgments
Supplementary Materials
@@
Simulated implosions are modeled using a radiation drive derived from the Callahan hohlraum model 7 and solve the rocket equations for 200 coupled rocket-pistons that discretize the capsule shell in 3D. After the initial acceleration phase, the pistons are linked through hotspot pressure using power-balance equations, following the framework of Springer et al. 38 but with an explicit hohlraum model applied from the onset of implosion. Synthetic diagnostics at stagnation and peak burn are generated through post-processing. Full methodological details are provided in 9.
@@
An ensemble of 443,610 RP simulations was used with inputs sampled near the predicted experimental ignition cliff 9. A random split of 986 samples (0.2%) was held out for testing, and 100 additional samples were used for validation during training. Images were normalized to by dividing by max pixel intensity for each view and channel. Scalar inputs and outputs were normalized to given mean and then passed through an inverse sigmoid to more closely approximate zero-centered Gaussians. All evaluation metrics reported in Tables 1 and 2 and Figure 3 are performed after removing normalization.
@@
DDPMs generate high resolution data distributions by first noising training data to gradually approach a Gaussian distribution, then iteratively denoising samples drawn from a Gaussian prior 15, 36. During training, samples can be drawn at arbitrary diffusion times according to , where , and corresponding to the noise schedule. We use a linear noise schedule that decreases from 0.9999 to 0.98 over 1000 steps.
@@
A multi-modal neural network architecture is used to encode and predict noise in both scalar and image pixel space. We adapt a U-net architecture 35 developed by 12 to ingest and predict multi-modal data. Each image view is embedded by independent convolutions networks containing 48 channels and 3 up/down sampling layers. Both input scalars, output scalars, and mask tokens are embedded by fully connected neural networks containing one 64-dimensional hidden layers. Diffusion time is embedded by a sin/cos positional encoding. Scalar, mask, and time embeddings are concatenated and added to U-net embedding at each up and down sampling step. Image embedding for each view receive the same time, scalar, and mask conditioning but are independent from each other until the bottleneck layer (lowest dimension of the U-net). At the bottleneck layer outputs from each image view are pooled along the channel dimension, flattened, and concatenated along with scalar and mask embedding. Output and input scalar decoders take the flattened bottleneck layer as input and predict noise for each scalar. Image decoders pool information from the bottleneck layer and upsample three times to recover the original pixel and channel dimensions.
@@
Binary masks inform the model if a conditioning variable (scalar or image) is provided for conditioning, or if it should be predicted during the denoising (inference) process. Certain masks are linked to others if those quantities can never be known without the other. Specifically, all inputs masks are linked since they must all be specified to perform a simulation and cannot be partially observed in experiment. The four inferred outputs scalars (R Weighted Harmonic Mean, R Mean, RKE at Bang-time, and RKE at Minimum Volume) are linked for the same reason 19. The three velocity scalars are linked because they are x,y,z components of the same vector. All other output scalars have individual masks along with each image view and channel. During training, all independent masks are sampled according to a Bernoulli distribution with constant 0.5. Ablations were performed with varied constant values and with curriculum learning strategies to gradually lower the constant as function of training time, but a fixed constant of 0.5 provided optimal performance across forward, inverse, and imputation tasks.